3(x-4)- (x+1) <,= -12


trying to figure out why the answer is x <,= 1/2



can anyone help

Answer by rfer(10417)

as you work the equation you get down to
3x-12-x-1<=-12
combine
2x-13<=-12
add 13 to both sides
2x-13+13<=-12+13
combine
2x=+1
divide by 2
2x/2<=1/2
x<=1/2